OpenStudy (iwanttogotostanford):

how exactly would i do this ? Determine if the following function is even, odd, or neither. f(x) = -9x^4 + 5x + 3

OpenStudy (iwanttogotostanford):

@welshfella @rock_mit182

OpenStudy (iwanttogotostanford):

I know i have to replace x with -x and see if it stays the same or not but I'm still confused

OpenStudy (freckles):

what did you get after replacing x with -x that is what is f(-x)=?

OpenStudy (iwanttogotostanford):

I got -9 (-x^4)+5-x+3

OpenStudy (freckles):

\[f(x)=-9x^4+5x+3\] \[f(-x)=-9(-x)^4+(-x)+3\]

OpenStudy (freckles):

do you know how to simplify (-x)^4?

OpenStudy (iwanttogotostanford):

no... how do you?

OpenStudy (freckles):

do you know what (-1)^4=?

OpenStudy (iwanttogotostanford):

yes, its just 1

OpenStudy (freckles):

so (-x)^4=?

OpenStudy (iwanttogotostanford):

oh just 1

OpenStudy (iwanttogotostanford):

or x

OpenStudy (freckles):

well no you have x inside the 4th power

OpenStudy (freckles):

(-x)^4 is the same as (-1*x)^4 which is the same as (-1)^4*x^4 by law of exponents and you just said (-1)^4 is 1 so (-x)^4=x^4

OpenStudy (iwanttogotostanford):

ok so now we have:f(−x)=−9(−x)4+(−x)+3

OpenStudy (freckles):

\[(-x)^{even number }=x^{even number} \\ (-x)^{odd number } =- (x^{odd number}) \text{ or we just write this as } -x^{oddnumber}\]

OpenStudy (freckles):

4 is even number so (-x)^4 = x^4

OpenStudy (iwanttogotostanford):

so its even right? since it stayed the same

OpenStudy (freckles):

we haven't got to the other terms yet

OpenStudy (iwanttogotostanford):

oh ok so what do we have so far just so i know whats going on exactly

OpenStudy (freckles):

are you understanding we have only played with the first term in f(-x)?

OpenStudy (freckles):

you do see that (-x)^4 is x^4 ?

OpenStudy (iwanttogotostanford):

yes i do understand

OpenStudy (freckles):

so that is as far as we have gotten so far :p

OpenStudy (iwanttogotostanford):

ok

OpenStudy (freckles):

\[f(x)=-9x^4+5x+3 \\ f(-x)=-9(-x)^{4}+5(-x)+3 \\ f(-x)=-9x^4-5x+3\]

OpenStudy (freckles):

so looking at f(x) and f(-x) do you think they are the same ?

OpenStudy (freckles):

compare the terms are all the terms the same?

OpenStudy (iwanttogotostanford):

no

OpenStudy (freckles):

the middle term is different correct?

OpenStudy (iwanttogotostanford):

yes

OpenStudy (freckles):

anyways same=even if you multiply f(-x) by -1 and get f(x) then is odd so if you multiply every term in f(-x) by -1 is the result the same as the terms in f(x) ?

OpenStudy (freckles):

like we already said the function f(x) wasn't even because you said f(x) and f(-x) weren't the same

OpenStudy (iwanttogotostanford):

if you multiply every term f(-x) by -1 it would not be the same as the other function

OpenStudy (freckles):

so f is not odd either

OpenStudy (freckles):

there is a short cut if you want to know

OpenStudy (freckles):

before I move on to the short cut do you understand the answer ?

OpenStudy (iwanttogotostanford):

yes i want to know the short cut

OpenStudy (iwanttogotostanford):

so it is neither? I don't quite understand that/....

OpenStudy (freckles):

well if it isn't even and if it isn't odd then it is neither odd or even

OpenStudy (iwanttogotostanford):

and why were there 3 functions when we were comparing since i thought there were only two?

OpenStudy (freckles):

f(x) wasn't the same as f(-x) so f(x) was not even f(x) wasn't the same as -f(-x) so f(x) was not odd

OpenStudy (freckles):

i'm using the definition of even and odd function

OpenStudy (iwanttogotostanford):

i know that but why were there three functions when comparing?

OpenStudy (freckles):

the definition of an even function must satisfy f(x)=f(-x) the definition of odd function must satisfy f(x)=-f(-x)

OpenStudy (iwanttogotostanford):

since there is only an even and odd

OpenStudy (freckles):

that is why we has to find f(-x) and also -f(-x)

OpenStudy (iwanttogotostanford):

oh so it was the given function, the odd function, and the even funciton>?

OpenStudy (freckles):

our function was neither odd or even

OpenStudy (iwanttogotostanford):

ok, i get that now. So what is the short cut to this?

OpenStudy (freckles):

ok do you know how to find the degree of each term of your polynomial ?

OpenStudy (freckles):

\[f(x)=-9x^4+5x+3\]

OpenStudy (iwanttogotostanford):

sort of, I'm a little bit fuzzy on that though i learned it last year

OpenStudy (freckles):

what is the degree of the first term? the degree of the second? the degree of the third?

OpenStudy (freckles):

hint look at the exponent on the variable

OpenStudy (iwanttogotostanford):

1st- 4, 2nd- 2, 3rd-1 ? I'm probably wrong there ...

OpenStudy (freckles):

ok this might help you our function is equivalent to \[f(x)=-9x^4+5x^{\color{red}{1}}+3\color{red}{x^{0}}\]

OpenStudy (freckles):

if you see a variable and no power the power is understood to be one

OpenStudy (iwanttogotostanford):

oh ok, that was the other option i was debating in my head. ok, got it.

OpenStudy (freckles):

if there is no variable then the degree is 0

OpenStudy (freckles):

so if you have a polynomial and all the powers are even then the function is even if you have a polynomial and all the powers are odd then the function is odd if you have a polynomial and you have a mixture of even and odds in your powers then the function is neither even or odd

OpenStudy (freckles):

the powers were 4,1,0 are all of these numbers even , odd, or do you have a mixture?

OpenStudy (iwanttogotostanford):

oh wow, thats a big help. Thank you so much for sharing the tip! And we have a mixture since 0 is even and 1 is odd

OpenStudy (freckles):

right

OpenStudy (freckles):

so the function f(x)=-9x^4+5x+3 is neither even or odd

OpenStudy (freckles):

what about f(x)=9x^4+3 ?

OpenStudy (iwanttogotostanford):

is would be a mix also since 3 would have degree of 0, and 4 is even

OpenStudy (freckles):

isn't 4 and 0 even?

OpenStudy (iwanttogotostanford):

oh yeah duh wow that was dumb. yes,i knew that. the function would be even.

OpenStudy (freckles):

yep

OpenStudy (freckles):

there is one exception

OpenStudy (freckles):

the function f(x)=0 is actually both even and odd

OpenStudy (iwanttogotostanford):

oh ok

OpenStudy (freckles):

but anyways it will work every other time for any other polynomial function

OpenStudy (iwanttogotostanford):

ok. thanks! That helped a lot. can you help me with another question I have if i tag you in a new post?

OpenStudy (freckles):

the reason f(x)=0 is both even and odd because it does satisfy both definitions f(-x)=0 so f(x)=f(-x) so f is even -f(-x)=-0=0 so f(x)=-f(-x) so f is odd

OpenStudy (freckles):

ok

OpenStudy (iwanttogotostanford):

ok!

OpenStudy (iwanttogotostanford):

thanks for the help. Ill tag you now.